OpenStudy (anonymous):

For PSET 1 series 1 (vectors and dot) products. 1A: question 6 asks to find the velocity vector a of a plane that travels 200 mph north despite a velocity vector of wind coming from the north east. My answer was {(25*sqrt(2)*i} + {(200+25*sqrt(2)*j}, however the solutions page gives the answer {(50/sqrt(2))*i} + {(200+(50/sqrt(2))*j}. Am I wrong or is the solution wrong. And if I am what am I doing wrong?

OpenStudy (anonymous):

\[\frac{ 50 }{ \sqrt(2) } = 25\sqrt(2)\]

OpenStudy (anonymous):

It is?

OpenStudy (anonymous):

I did it on the calculator. Thank you for your help.